# Collective dynamics versus entanglement in quantum battery performance

### v2

@@ IV Conclusion
In this work, we have investigated the physical mechanisms underlying energy storage and power enhancement in many-body quantum batteries, with particular emphasis on the role of quantum correlation and interaction structure, and the number of particles involved in entanglement. By systematically comparing energetic quantities such as stored energy and instantaneous charging power with a hierarchy of information-theoretic measures, we have clarified how entanglement and many-body dynamics jointly shape the charging process.
A central finding of our study is the robust temporal separation between energy transfer and correlation buildup. Across all battery–charger configurations considered, the instantaneous power consistently reaches its maximum before bipartite, tripartite, and multipartite entanglement attain their peak values. This establishes that peak charging is governed primarily by coherent collective transport enabled by interactions, rather than by fully developed entanglement or scrambling. Quantum correlations, while essential for stabilizing and redistributing energy at later times, do not directly determine the timing of maximal power output. This resolves an important ambiguity in the literature regarding the role of entanglement as a driver versus a byproduct of efficient charging.
Extending our analysis to $\kappa$-local charging protocols allowed us to disentangle genuine collective quantum effects from trivial classical scaling. Under unconstrained conditions, enhanced charging can be fully attributed to increases in the energy scale of the charger Hamiltonian. Imposing a fixed operator norm, thereby enforcing fair charging reveals that increasing the interaction order or the number of participating particles does not automatically lead to higher charging power. Charging performance is determined by the number of particles that are actually mutually correlated and contribute to entanglement. Fully collective interactions offer a genuine advantage, as all particles participate coherently in the charging process. In contrast, partially extended interaction schemes do not consistently increase the number of effectively interacting particles, and therefore, cannot guarantee improved charging efficiency.
Extending this analysis, adding NNN couplings in NN coupling under fair charging conditions further illustrates the limitations of partial interaction extensions. Compared to NN interactions alone, the maxima of instantaneous power are suppressed, confirming that merely increasing the participation number does not enhance charging performance.

### v3

@@ IV Conclusion
In this work, we investigated the physical mechanisms underlying energy storage and power enhancement in many-body quantum batteries, with a focus on the role of quantum correlations, interaction structure, and the extent of the multipartite nature of entanglement. By systematically comparing energetic quantities such as stored energy and instantaneous charging power with a hierarchy of entanglement measures, we tried to clarify how entanglement and many-body dynamics jointly govern the charging process.
We began with two-local interacting systems and observed a robust temporal separation between energy transfer and correlation buildup. Across all battery–chargoid configurations considered, the instantaneous power consistently reaches its maximum before bipartite, tripartite, and multipartite entanglement attain their peak values, indicating that coherent collective dynamics is the dominating factor for peak charging.
Extending the analysis to $\kappa$-local charging protocols, we considered both unconstrained and constrained forms of the chargoid Hamiltonian. Under unconstrained conditions, enhanced charging arises primarily from the increased energy scale of the chargoid Hamiltonian. By imposing a constraining condition, viz. fixed operator norm (“fair” charging), we found that increasing the interaction order or the number of participating particles does not automatically improve charging power. Instead, performance is dictated by the number of particles that are actually mutually correlated and contribute to entanglement. Fully collective interactions provide a genuine advantage, as all particles participate coherently in the charging process. In contrast, partially extended interaction schemes fail to involve all particles, and do not reliably enhance charging efficiency.
Further, incorporating NNN interactions along with NN couplings under fair charging conditions illustrated the limitations of partial interaction extensions. Compared to NN interactions alone, the maxima of instantaneous power are suppressed, confirming that merely increasing the nominal participation number without achieving full collective correlations is insufficient to enhance charging performance.